(px-y)( x-py)=2p
Using x^2=u,y^2=u
Find general solution and solveble solutin
1
Expert's answer
2020-10-21T17:23:31-0400
"(px-y)(x-py)=2p"

"(pxy-y^2)(x-py)=2py"

"({py \over x}\cdot x^2-y^2)(1-{py \over x})=2{py \over x}"

"{py \over x}\cdot x^2-y^2=\dfrac{2\dfrac{py}{x}}{1-\dfrac{py}{x}}"

"y^2={py \over x}\cdot x^2-\dfrac{2\dfrac{py}{x}}{1-\dfrac{py}{x}}"

Let "x^2=u, y^2=v." Then


"du=2xdx, dv=2ydy"

"p=\dfrac{dy}{dx}=\dfrac{2x}{2y}\cdot \dfrac{dv}{du}=\dfrac{\sqrt{u}}{\sqrt{v}}\cdot P, P=\dfrac{dv}{du}"

"{py \over x}=\dfrac{dv}{du}=P"

Substitute


"v=Pu-\dfrac{2P}{1-P}"

We have Clairaut Equation


"v=Pu+f(P)"

The general solution is given by "v=Cu+f(C), C" is an arbitrary constant.

Therefore


"v=Cu-\dfrac{2C}{1-C}, C \ \text{is arbitrary constant}, C\not=1"

The general solution is


"y^2=Cx^2-\dfrac{2C}{1-C}, \ C \ \text{is arbitrary constant}, C\not=1"